OpenStudy (anonymous):

Solve the system algebraically. Check your work. 4x + 2y = 10 3x + 2y = 6 Make sure there are NO SPACES in your answer. Be sure to include a comma. ANSWER: {( ________a0 )}

OpenStudy (amistre64):

let x= (5+n)/4 let y = (5-n)/2

OpenStudy (anonymous):

k